OpenStudy (anonymous):

help on 4, 6, 8 Find the values of X and Y for which ABCD must be a parallelagram.

OpenStudy (anonymous):

Thanks

OpenStudy (mertsj):

3x+1=6x-4

Directrix (directrix):

4. If one pair of opposite sides of a quadrilateral is both parallel and congruent, then the quadrilateral is a parallelogram. So, 3x+1 = 6x-4 and x = 5/3.

OpenStudy (mertsj):

4x-1=x+38 (if I saw the numbers right. They are very small)

OpenStudy (anonymous):

so then X and Y would be for number 4?

OpenStudy (mertsj):

Number 8 I don't think so but I can't see it very well. only thing I can see is that one diagonal is bisected.

OpenStudy (anonymous):

ill re send it

OpenStudy (anonymous):

Determine weather the quadratalerl is a prallelagram

Directrix (directrix):

x = 5/3. There is no y variable. See original post above.

OpenStudy (anonymous):

Oh I see and for #6 would it be 13?

Directrix (directrix):

Agree with Mert on #8. The quad is not a parallelogram.

OpenStudy (anonymous):

what about #

OpenStudy (anonymous):

6

Directrix (directrix):

x =13 on #6 because one pair of opposite sides is both congruent and parallel.